# GNSS-based Lunar Orbit and Clock Estimation With Stochastic Cloning UD Filter

### v1

@@ 5.5 Ionospheric and Plasmapsheric Delays
The ionospheric and plasmaspheric delays are modeled by integrating the rays along the GNSS-to-lunar satellite. The path of the ray is governed by the refractive index (which can be computed from electron density) profile of the ionosphere and plasmasphere, which are simulated using the International Reference Ionosphere (IRI) model [Bilitza2022IRIReview] for the ionosphere and the Global Core Plasma Model (GCPM) [Gallagher2000GCPM] for the plasmasphere. Figure 7 shows the simulated electron density profiles at the equator and 45 degrees latitude at March 1, 2025, 12:00:00 UTC, which is the initial epoch of the simulation.
@@
Figure 7: The simulated electron density profiles at March 1, 2025, 12:00:00 UTC using the IRI and GCPM models at the X-Y plane (left) and the X-Z plane (right) in the solar magnetic coordinate system (X:-X: toward the Sun, $Z:$ Earth’s geomagnetic pole). The $K_{p}$ index is set to 3, and the Rz12 value is set to 50. The color bar indicates the electron density in units of electrons per cubic centimeter. The blue, orange, and green lines represent the orbit altitudes of the GPS (20,200 km), Galileo (23,222 km), and GEO (35,786 km), respectively.
@@ 6.1 Satellite Availability
Figure 8 shows the satellite availability for a lunar GNSS receiver over the entire simulation period. A satellite is considered available if the receiver can track the satellite signal with a carrier-to-noise density ratio (C/N0 ) greater than 18 dB-Hz. Note that [Delepaut2020GNSSLunar] points out that higher C/N0 will be required to decode the almanac and ephemeris from the navigation messages. For this paper, we assume the ephemeris information is provided by uplink from the ground.
@@
Note that since the Earth’s direction is nearly orthogonal to the orbital plane of the lunar satellite, we do not observe significant outages in satellite availability during the simulation period, due to lunar occultation of the GNSS satellites. In future works, we will investigate different lunar satellite orbits with different right ascension of ascending node (RAAN) to analyze the effect of lunar occultation on the estimation accuracy.
@@ 6.2 Satellite Ephemeris Errors
The ephemeris error distributions for the GPS (L1/L5), Galileo (E1/E5a), and QZSS (L1/L5) satellites in directions along the GNSS-to-Lunar satellite line-of-sight (LOS) is shown in Fig. 9. We see that most of the position and clock bias ephemeris errors are within $\pm 2$ m. The total user range error from the ephemeris are mostly within $\pm 4$ m, and has different mean biases for different GNSS constellations. The mean and variances of the ephemeris errors in LOS direction for different GNSS constellations are summarized in Table 6.
@@ 6.6 Summary
Pseudorange measurements offer the highest availability but are strongly affected by ionospheric and plasmaspheric delays, particularly at low-to-medium tangential altitudes. Using single-frequency pseudorange without introducing significant biases requires one (or more) of the following: (i) aggressive tangential-altitude screening, (ii) inflating measurement variances (suboptimal but simple), (iii) compensation using external delay models, or (iv) explicit estimation (or addition as ‘consider’ states) of delay states in the filter. Ionosphere-free pseudorange mitigates the dominant first-order delay but reduces availability and increases receiver noise. TDCP provides a favorable compromise when tangential-altitude screening is applied, because differencing removes integer ambiguity and suppresses slowly varying biases while retaining mm-level precision when the ionospheric delays are small compared to receiver noise.
@@ 7.1 Filter Settings
We evaluate the proposed UDU filter and fixed-interval smoother under three measurement configurations: (1) L1 pseudorange only, (2) ionosphere-free (IF) pseudorange only, and (3) IF pseudorange combined with TDCP. Configuration (1) serves as a baseline and follows common assumptions in prior lunar GNSS ODTS studies [Iiyama2024TDCPNavigation, Mina2025LCRNS]: we process L1 pseudorange with a tangential-altitude mask of 1000 km to reject rays that traverse the ionosphere. Configuration (2) replaces single-frequency pseudorange with IF pseudorange using the same 1000 km mask to further suppress higher-order ionospheric terms and bending effects that are not removed by the first-order ionosphere-free combination. Configuration (3) combines IF pseudorange (1000 km mask) with TDCP measurements, applying a stricter mask of 5000 km for TDCP to preferentially retain measurements for which time-differenced ionospheric/plasmaspheric residuals are small relative to carrier-phase noise.
@@
For each measurement configuration, we run 50 Monte-Carlo simulations with independent realizations of the initial state perturbations and measurement errors. To ensure convergence from the large initial uncertainties, performance is evaluated over the final orbit revolution (approximately 30 hours). We report signal-in-space error (SISE) metrics consistent with the LunaNet SRD [nasa2022srd]:
@@ 7.2 Results
Table 9: SISE position and velocity errors of the filter and smoother under different measurement configurations. Values are the means of over 50 Monte-Carlo simulations. (PR=pseudorange, TDCP=time-differenced carrier phase, IF=ionosphere-free).
@@ 7.2.1 L1 Pseudorange Only (Baseline)
Using only L1 pseudorange yields the largest SISE errors for both position and velocity. This is consistent with unmodeled plasmaspheric delays contaminating pseudorange even after applying the 1000 km tangential-altitude mask. These results suggest that pseudorange-only filtering requires either (i) more aggressive variance inflation to maintain consistency or, more optimally, (ii) explicit modeling/estimation of additional bias states (e.g., plasmaspheric delay and LOS range biases) to prevent systematic errors. For single-frequency receivers, an alternative is the GRAPHIC observable [Bock2009GPS_SF_Orbit, Shi2012IonosphericPPP], which suppresses first-order plasmaspheric delay by combining code and phase; however, this requires estimating float ambiguities, increasing state dimension and implementation complexity.
@@ 7.2.2 Ionosphere-free Pseudorange Only
Replacing single-frequency pseudorange with ionosphere-free pseudorange substantially improves both position and velocity SISE, despite reduced availability (dual-frequency tracking is required, and fewer satellites transmit both signals). This improvement indicates that first-order plasmaspheric delays are a dominant error source for the baseline configuration, and that IF measurements provide a practical mitigation in the absence of an explicit plasmaspheric delay estimator.
@@ 7.2.3 Ionosphere-free Pseudorange + TDCP
Adding TDCP further improves SISE performance for both the filter and smoother. TDCP provides high-precision relative information between successive epochs, tightening constraints on velocity- and clock-drift-related states and reducing overall estimation error. Relative to IF pseudorange alone, the filter RMS SISE improves by approximately 16–18% for position and 9–11% for velocity.
@@ 7.2.4 Effect of Smoothing
Applying the fixed-interval smoother improves both position and velocity SISE for all measurement configurations. The TDCP-assisted smoother achieves the best overall performance. Figure 14 compares the filter and smoother trajectories (after five filter–smoother iterations) for one Monte-Carlo run with IF pseudorange and TDCP. Large position SISE errors near apolune, where slower orbital motion reduces instantaneous observability, and larger velocity SISE errors near perilune are both noticeably reduced after smoothing.
Smoothing also suppresses the occasional filter spikes induced by near-outlier TDCP updates, reducing the 99.7% velocity SISE from $1.90\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ to $0.80\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ (58% reduction). Overall, the smoothed velocity SISE with TDCP is approximately 30% lower than the IF pseudorange-only smoother.
@@ 7.2.5 Estimation Error Distribution
These trends are also reflected in the SISE histograms over the final revolution across 50 Monte-Carlo runs (Figs. 15 and 16). Incorporating TDCP shifts both the position and velocity SISE distributions toward smaller values for both filtering and smoothing. Comparing filter versus smoother results shows that smoothing reduces the tails of the SISE distributions, consistent with its ability to incorporate future measurements to refine poorly observed portions of the orbit.
Figure 15: Histogram of filter position and velocity SISE over 50 Monte-Carlo simulations (final revolution only).
Figure 16: Histogram of smoother position and velocity SISE over 50 Monte-Carlo simulations (final revolution only).
@@ 8 Conclusion
This paper presented a terrestrial GNSS–based orbit and clock estimation framework for lunar navigation satellites that combines stochastic cloning with a numerically robust UD-factorized filter and a delayed-state fixed-interval smoother. To address the low observability and numerical sensitivity that arise when incorporating precise time-differenced carrier phase (TDCP) measurements at lunar distances, we developed a stable covariance-factorization pipeline that can process delayed-state measurements without sacrificing positive semi-definiteness. In addition, we formulated dynamics and measurement models that explicitly include relativistic coupling between orbit and clock states, transformations among Earth and lunar time scales, and propagation delays through the terrestrial ionosphere and plasmasphere.
We evaluated the proposed approach using high-fidelity Monte-Carlo simulations that incorporate realistic multi-constellation GNSS geometry, broadcast-ephemeris errors, and ray-tracing-based ionospheric/plasmaspheric delay models. The results show that augmenting ionosphere-free pseudorange with TDCP consistently improves estimation performance relative to pseudorange-only baselines. In particular, TDCP reduces the filtering RMS position and velocity SISE by approximately 16–18% and 9–11%, respectively, achieving $3.68\text{\,}\mathrm{m}$ RMS position SISE and $0.39\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ RMS velocity SISE under the tested configuration. The delayed-state smoother further reduces residual errors and tails in the error distribution, yielding sub-$4\text{\,}\mathrm{m}$ position accuracy and sub-$\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ velocity accuracy, which satisfies the stringent accuracy targets envisioned for future Lunar Augmented Navigation Services (LANS).
Future work will focus on (i) systematic tuning of tangential-altitude masking and measurement noise inflation to better trade measurement availability against plasmaspheric-delay biases, and (ii) reducing reliance on masking by explicitly estimating residual ionospheric/plasmaspheric terms and multi-constellation inter-system biases within the state (or as ‘consider’ parameters). We will also investigate processing raw pseudorange and GRAPHIC observables with ambiguity handling to improve robustness and availability for single-frequency or intermittently dual-frequency reception scenarios.

### v2

@@ 5.5 Ionospheric and Plasmapsheric Delays
The ionospheric and plasmaspheric delays are modeled by integrating the rays along the GNSS-to-lunar satellite. The path of the ray is governed by the refractive index (which can be computed from electron density) profile of the ionosphere and plasmasphere. The simulation uses the International Reference Ionosphere (IRI) model [Bilitza2022IRIReview] for the ionosphere and the Global Core Plasma Model (GCPM) [Gallagher2000GCPM] for the plasmasphere. Figure 8 shows the simulated electron density profiles at the equator and 45 degrees latitude at March 1, 2025, 12:00:00 UTC, which is the initial epoch of the simulation. To model the electron densities under moderate solar and geomagnetic activity, we set the $K_{p}$ index is set to 3, and the 12-month running mean of the daily sunspot number (Rz12) to 50 for this plot and the rest of the paper.
@@
Figure 8: The simulated electron density profiles at March 1, 2025, 12:00:00 UTC using the IRI and GCPM models at the X-Y plane (left) and the X-Z plane (right) in the solar magnetic coordinate system (X:-X: toward the Sun, $Z:$ Earth’s geomagnetic pole). The color bar indicates the electron density in units of electrons per cubic centimeter. The blue, orange, and green lines represent the orbit altitudes of the GPS (20,200 km), Galileo (23,222 km), and GEO (35,786 km), respectively.
@@ 6.1 Satellite Availability
Figure 9 shows the satellite availability for a lunar GNSS receiver over the entire simulation period. A satellite is considered available if the receiver can track the satellite signal with a C/N0 greater than 18 dB-Hz. Note that [Delepaut2020GNSSLunar] points out that higher C/N0 will be required to decode the almanac and ephemeris from the navigation messages. For this paper, we assume the ephemeris information is provided by uplink from the ground.
@@
Note that since the Earth’s direction is nearly orthogonal to the orbital plane of the lunar satellite, we do not observe significant outages in satellite availability during the simulation period, due to lunar occultation of the GNSS satellites. In future works, we will investigate different lunar satellite orbits with different right ascension of ascending nodes (RAAN) to analyze the effect of lunar occultation on the estimation accuracy.
@@ 6.2 Satellite Ephemeris Errors
The ephemeris error distributions for the GPS (L1/L5), Galileo (E1/E5a), and QZSS (L1/L5) satellites in directions along the GNSS-to-Lunar satellite line-of-sight is shown in Fig. 10. We see that most of the position and clock bias ephemeris errors are within $\pm 2$ m. The total user range error from the ephemeris are mostly within $\pm 4$ m, and has different mean biases for different GNSS constellations. The mean and variances of the ephemeris errors in LOS direction for different GNSS constellations are summarized in Table 6.
@@ 6.6 Summary
Pseudorange measurements offer the highest availability but are strongly affected by ionospheric and plasmaspheric delays, particularly at low-to-medium tangential altitudes. Using single-frequency pseudorange without introducing significant biases requires one (or more) of the following: (i) aggressive tangential-altitude screening, (ii) inflating measurement variances (suboptimal but simple), (iii) compensation using external delay models, or (iv) explicit estimation (or addition as consider states where cross-covariance is taken into account but the states are not explicitly estimated [Tapley2004]) of delay states in the filter. Ionosphere-free pseudorange mitigates the dominant first-order delay but reduces availability and increases receiver noise. TDCP provides a favorable compromise when tangential-altitude screening is applied, because differencing removes integer ambiguity and suppresses slowly varying biases while retaining mm-level precision when the ionospheric delays are small compared to receiver noise.
@@ 7.1 Filter Settings
We evaluate the proposed UDU filter and fixed-interval smoother under three measurement configurations: (1) L1 pseudorange only, (2) ionosphere-free pseudorange only, and (3) ionosphere-free pseudorange combined with TDCP. Configuration (1) serves as a baseline and follows common assumptions in prior lunar GNSS ODTS studies [Iiyama2024TDCPNavigation, Mina2025LCRNS]: we process L1 pseudorange with a tangential-altitude mask of 1000 km to reject rays that traverse the ionosphere. Configuration (2) replaces single-frequency pseudorange with ionosphere-free pseudorange using the same 1000 km mask to further suppress higher-order ionospheric terms and bending effects that are not removed by the first-order ionosphere-free combination. Configuration (3) combines ionosphere-free pseudorange (1000 km mask) with TDCP measurements, applying a stricter mask of 5000 km for TDCP to preferentially retain measurements for which time-differenced ionospheric/plasmaspheric residuals are small relative to carrier-phase noise.
@@
For each measurement configuration, we run 50 Monte-Carlo simulations with independent realizations of the initial state perturbations and measurement errors. To ensure convergence from the large initial uncertainties, performance is evaluated over the final orbit revolution (approximately 30 hours). We report SISE metrics consistent with the LunaNet SRD [nasa2022srd] as follows:
@@ 7.2 Results
Table 9: SISE position and velocity errors of the filter and smoother at the final orbit under different measurement configurations. Values are the means of over 50 Monte-Carlo simulations. (PR=pseudorange, TDCP=time-differenced carrier phase, IF=ionosphere-free).
@@ 7.2.1 L1 Pseudorange Only (Baseline)
Using only L1 pseudorange yields the largest SISE errors for both position and velocity. This is consistent with unmodeled plasmaspheric delays contaminating pseudorange even after applying the 1000 km tangential-altitude mask. These results suggest that pseudorange-only filtering requires either (i) more aggressive variance inflation to maintain consistency or, more optimally, (ii) explicit modeling/estimation of additional bias states (e.g., plasmaspheric delay and LOS range biases) to prevent systematic errors. For single-frequency receivers, an alternative is the Group and Phase Ionosphere Correction (GRAPHIC) observable [Bock2009GPS_SF_Orbit, Shi2012IonosphericPPP], which suppresses first-order plasmaspheric delay by combining code and phase; however, this requires estimating float ambiguities, increasing state dimension and implementation complexity.
@@ 7.2.2 Ionosphere-free Pseudorange Only
Replacing single-frequency pseudorange with ionosphere-free pseudorange substantially improves both position and velocity SISE, despite reduced availability (dual-frequency tracking is required, and fewer satellites transmit both signals). This improvement indicates that first-order plasmaspheric delays are a dominant error source for the baseline configuration, and that ionosphere-free measurements provide a practical mitigation in the absence of an explicit plasmaspheric delay estimator.
@@ 7.2.3 Ionosphere-free Pseudorange + TDCP
Adding TDCP further improves SISE performance for both the filter and smoother. TDCP provides high-precision relative information between successive epochs, tightening constraints on velocity- and clock-drift-related states and reducing overall estimation error. Relative to ionosphere-free pseudorange alone, the filter RMS SISE improves by approximately 16–18% for position and 9–11% for velocity.
@@ 7.2.4 Effect of Smoothing
Applying the fixed-interval smoother improves both position and velocity SISE for all measurement configurations. The TDCP-assisted smoother achieves the best overall performance. Figure 15 compares the filter and smoother trajectories (after five filter–smoother iterations) for one Monte-Carlo run with ionosphere-free pseudorange and TDCP. Large position SISE errors near apolune, where slower orbital motion reduces instantaneous observability, and larger velocity SISE errors near perilune are both noticeably reduced after smoothing.
Smoothing also suppresses the occasional filter spikes induced by near-outlier TDCP updates, reducing the 99.7% velocity SISE from $1.90\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ to $0.80\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ (58% reduction). Overall, the smoothed velocity SISE with TDCP is approximately 30% lower than the ionosphere-free pseudorange-only smoother.
@@ 7.2.5 Estimation Error Distribution
These trends are also reflected in the SISE histograms over the final orbit across 50 Monte-Carlo runs (Figs. 16 and 17). Incorporating TDCP shifts both the position and velocity SISE distributions toward smaller values for both filtering and smoothing. Comparison of filter versus smoother results shows that smoothing reduces the tails of the SISE distributions, consistent with its ability to incorporate future measurements to refine poorly observed portions of the orbit.
Figure 16: Histogram of filter position and velocity SISE over 50 Monte-Carlo simulations (final orbit only).
Figure 17: Histogram of smoother position and velocity SISE over 50 Monte-Carlo simulations (final orbit only).
@@ 8 Conclusion
This paper presented a terrestrial GNSS–based orbit and clock estimation framework for lunar navigation satellites that combines stochastic cloning with a numerically robust UD-factorized filter and a delayed-state fixed-interval smoother. To address the low observability and numerical sensitivity that arise when incorporating precise TDCP measurements at lunar distances, we developed UD-factorization filter that can process delayed-state measurements with stochastic cloning. In addition, we formulated dynamics and measurement models that explicitly include relativistic coupling between orbit and clock states, transformations among Earth and lunar time scales, and propagation delays through the terrestrial ionosphere and plasmasphere.
We evaluated the proposed approach using high-fidelity Monte-Carlo simulations that incorporate realistic multi-constellation GNSS geometry, broadcast-ephemeris errors, and ray-tracing-based ionospheric/plasmaspheric delay models. The results show that augmenting ionosphere-free pseudorange with TDCP consistently improves estimation performance relative to pseudorange-only baselines. In particular, TDCP reduces the filtering RMS position and velocity SISE by approximately 16–18% and 9–11%, respectively, achieving $3.68\text{\,}\mathrm{m}$ RMS position SISE and $0.39\text{\,}\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ RMS velocity SISE under the tested configuration. The delayed-state smoother further reduces residual errors and tails in the error distribution, yielding sub-$4\text{\,}\mathrm{m}$ position accuracy and sub-$\mathrm{m}\mathrm{m}\mathrm{/}\mathrm{s}$ velocity accuracy, which satisfies the stringent accuracy targets envisioned for future LANS.
Future work will focus on (i) systematic tuning of tangential-altitude masking and measurement noise inflation to better trade measurement availability against plasmaspheric-delay biases, and (ii) reducing reliance on masking by explicitly estimating residual ionospheric/plasmaspheric terms and multi-constellation inter-system biases within the state (or as consider states). We will also investigate processing raw pseudorange and GRAPHIC observables with ambiguity handling to improve robustness and availability for single-frequency or intermittently dual-frequency reception scenarios.